Summary Status

Siderostat
Distance
from
center (m)
SWS
EE
ME
OP
Last Auto-Collimation
Pointing Offsets to try
Remarks






North Periscope
South Periscope
North Periscope
South Periscope

N4
52.5
G
G
G
G
12/12/2012
12/12/2012

0, 0

N3
12.5
G
G
G
G
12/12/2012
2/7/2014
50, 0
-31, -32

N1
2.5
G
B
G
G
12/12/2012
12/12/2012
0, 0
-42, -39
See April 2013 trip report - apparently there is a dodgy Centronics connector.
S1
2.5
G
G
G
G
14/12/2012
14/12/2012
200, 0
75, 20

S2
7.5
G
G
G
G
12/12/2012
2/7/2014

0, 0

S3
27.5
G
G
G
G
14/12/2012
14/12/2012
120, -50
0, -100

S4
107.5
G
G
G
G
12/12/2012
12/12/2012


Seems OK; has behaved strangely in the past.
G=Good, B=Bad, Q=Questionable, SW=Software, EE=Electronics, ME=Mechanical, OP=Optics, A/C=auto-collimation

Baseline
Length (m)
Fringes (um)


North Periscope
South Periscope
N1-S1
5

-660
N1-S2
10

-800
N1-S3
30

-440
N1-S4
110


N3-S1
15

-700
N3-S2
20

-1770
N3-S3
40
?
-500
N3-S4
120


N4-S1
55


N4-S2
60
-400

N4-S3
80


N4-S4
160




Detailed Status


All siderostats seem occasionally to have trouble with azimuth indexing. Usually, they will find the index pulse within a couple of degrees if they miss the first one. There is some evidence that the issue is timing related (i.e. software is not checking often enough). There is also some evidence that it is related to the power supply (i.e. Vcc).

N1 - BAD

Occasionally working after a supplemental power supply was added in 2012.
Cannot be indexed properly, both azimuth and elevation axis have add issues, in 2014 only Azimuth is a problem.

Procedure after a power outage:
1) Auto-collimate with the green laser.
2) Do an "azindx" - You should be at 179.00.

If the star is out with this procedure, then adjust the azoff in sidstats.

N3 - OK

No comment.

N4 - OK

No comment.

S1 - OK

Periscope LEDs do not indicate on the rack in the control room when periscopes are out. They do indicate when periscopes are in.

S2 - OK

No comment.

S3 -OK

No comment.

S4 - OK?

Apparently, functioning correctly; has Behaved strangely in the past.
South periscope LEDs do not indicate on the rack in the control room.

Rest of SUSI Status


The RED laser seems to be non-functional. Is this because of the laser itself, the shutter or lots of ND? Who knows...

Last update: 3/7/2014